OpenStudy (anonymous):

Greg is in a car at the top of a roller-coaster ride. The distance, d, of the car from the ground as the car descends is determined by the equation d = 144 – 16t2, where t is the number of seconds it takes the car to travel down to each point on the ride. For which interval of time is Greg’s car moving in the air?

OpenStudy (anonymous):

\[0<t \le3\] \[-3<t \le3\] t>3 \[0<t<3\]

OpenStudy (error1603):

For the algebraic solution, you have to understand what d represents, Nora. "d" represents the distance the car is from the ground, so using the equation you are given: d=144-16t^2 you can substitute 0 for d because the car is going to be on the ground for your answer. then you solve for t by first subtracting 16t^2 from both sides 144=16t^2 than dividing both sides by 16 9=t^2 last take the square root both sides t=3